What is the specific heat of metal if its mass 27.32 g and it requires 108.3 J of heat energy to raise its temperature from 22.2⁰C to 32.5⁰C?
1
Expert's answer
2019-11-12T12:01:29-0500

Q=cm∆T.

c=Q/m∆T= 108.3/27.32 x 10.3= 0.384866878 J/g.